OpenStudy (anonymous):

Condense 1/2 log5 16 - 3 log5 X + 4 log5 Y

OpenStudy (shubhamsrg):

use these: 1) n log m = log(m^n) 2)log m + log n =log(mn) 3) log m -log n =log(m/n) will this be helpful ? note that base remains same in all the cases..(here it is 5)

OpenStudy (anonymous):

log5 (4y^4)/x^3

OpenStudy (anonymous):

thanks guys